OpenStudy (anonymous):

If f is a continuous function such that ∫^x f(t)dt=(3x)/4x^(2)+5) 0 find the value of f(1)

hartnn (hartnn):

Differentiate both sides w.r.t x so that you can use, \(\huge \frac{d}{dx} \int \limits_0^x f(t)dt=f(x)\) hence you get f(x) by differentiating. then just put x=1 in it.

hartnn (hartnn):

u understand what u need to do ? @paitucker

OpenStudy (anonymous):

so after you differentiate and get 3/8x what do i do?

hartnn (hartnn):

oh, there you need quotient (u/v) rule, you cannot just diff. numerator and denominator separately.

hartnn (hartnn):

\(\huge \frac{3x}{4x^2+5}\) is the question, right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

know how to use quotient rule ?

OpenStudy (anonymous):

yeah,im trying right now

OpenStudy (anonymous):

(4x^2+5)(3)-24x^2/16x^2+25 ?

hartnn (hartnn):

the denominator is incorrect

hartnn (hartnn):

it should be (4x^2+5)^2 don't simplify it.

hartnn (hartnn):

numerator is correct :)

OpenStudy (anonymous):

thanks :) , what should I do after that?

hartnn (hartnn):

so u have f(x) = ..... to find f(1), just put x=1 in that.

hartnn (hartnn):

f(x)=[(4x^2+5)(3)-24x^2]/(4x^2+5)^2 put x=1

OpenStudy (anonymous):

5/81 ?

hartnn (hartnn):

numerator = (4x^2+5)(3)-24x^2 = (4+5)*3-24 = 9*3 -24 = 27-24 = 3 denominator = 81 is correct.

hartnn (hartnn):

got your error ?

hartnn (hartnn):

so its 3/81 = 1/27

OpenStudy (anonymous):

ohh okay :) I get, thanks!!

hartnn (hartnn):

welcome ^_^